f(x+y)=f(x)f(y)
and f(1)=1/2
 
what’s the result of f(0)+f(1)+f(2)

Grade:12

2 Answers

Arun
4 years ago
Dear student
 
when
f(x+y)=f(x)f(y)
 
then f(x) = ax
now f(1) = 1/2
hence f(x) = (1/2)x
 
hence f(0) + f(1) + f(2) = 1 + ½ + ¼ = 7/4
Vikas TU
4 years ago
Dear student 
The question is incorrect , Please give the value of function , 
You can also upload the image . 
If we consider te value of function by own the ans would be fferent 
Good Luck